# Feature learning as alignment: a structural property of gradient descent in non-linear neural networks

### v3

@@ 7 Acknowledgements
We thank Lechao Xiao for detailed feedback on the manuscript. We also thank Jeffrey Pennington for helpful discussions. This work used the programs (1) XSEDE (Extreme science and engineering discovery environment) which is supported by NSF grant numbers ACI-1548562, and (2) ACCESS (Advanced cyberinfrastructure coordination ecosystem: services & support) which is supported by NSF grants numbers #2138259, #2138286, #2138307, #2137603, and #2138296. Specifically, we used the NCSA Delta system, via allocations TG-CIS220009.
@@ Appendix A Glossary
    $\bar{G}_{\ell}\equiv\frac{1}{n}\sum_{\alpha=1}^{n}\frac{\partial f(x^{(\alpha)})}{\partial x_{\ell}}\frac{\partial f(x^{(\alpha)})}{\partial x_{\ell}}^{\top}.$
@@ Appendix B Connection of the PTK to the empirical NTK
Appendix C Additional centerings of the NFC
@@
Appendix D Extending our theoretical predictions to depth and general activations
@@
Appendix E Free probability calculations of C-NFC
@@
We assume three properties hold in the finite dimensional case we consider, that will hold asymptotically in the infinite dimensional limit.
@@
Assumption 7 (Commutativity of expectation).
We will also make the approximation that the expectation commutes with ratio and square root.
$\mathbb{E}\left[\rho(A,B)\right]=\mathbb{E}\left[\frac{{\rm tr}(A^{\top}B)}{\sqrt{{\rm tr}(A^{\top}A){\rm tr}(B^{\top}B)}}\right]\approx\frac{\mathbb{E}\left[{\rm tr}(A^{\top}B)\right]}{\sqrt{\mathbb{E}\left[{\rm tr}(A^{\top}A)]{\rm E}[{\rm tr}(B^{\top}B)\right]}}$ (9)
@@
E.1 Free probability identities
@@ E.2 Numerator term of NFC
E.3 First denominator term of NFC
@@
E.4 Second denominator term of NFC
@@
Appendix F Alignment reversing dataset
The data consists of a mixture of two distributions from which two subsets of the data $X^{(1)}$ and $X^{(2)}$ are sampled from, and is parametrized by a balance parameter $\gamma\in(0,1]$ and two variance parameters $\epsilon_{1},\epsilon_{2}>0$. The subset $X^{(1)}$ which has label $y_{1}=1$ and constitutes a $\gamma$ fraction of the entire dataset, is sampled from a multivariate Gaussian distribution with mean 00 and covariance
$\Sigma=\bm{1}\bm{1}^{\top}+\epsilon_{1}\cdot I~{}.$
We construct $X^{(2)}$ by the following procedure:
@@ Appendix F Alignment reversing dataset
$\mathbb{E}\left[K^{2}_{ij}\right]$ $=\sum_{s}^{k}\sum_{\alpha}^{n}\sum_{p_{1},p_{2}}^{d}\mathbb{E}\left[a_{s}^{4}W_{s,p_{1}}X_{\alpha,p_{1}}X_{\alpha,p_{2}}W_{s,p_{2}}X_{i}^{\top}W_{s}W_{s}^{\top}X_{j}\right]$
$=\sum_{s}^{k}\sum_{\alpha}^{n}\sum_{p_{1},p_{2},q_{1},q_{2}}^{d}$
@@
Appendix G Varying the data distribution
@@
Appendix H Effect of initialization on feature learning
@@
Appendix I Experimental details
@@
Appendix J Additional SLO experiments
@@
Appendix K Additional C-NFC/UC-NFC plots across architectures
@@
Appendix L Additional experiments on real datasets

### v4

@@ 7 Acknowledgements
We thank Lechao Xiao for detailed feedback on the manuscript. We also thank Jeffrey Pennington for helpful discussions. This work used the programs (1) XSEDE (Extreme science and engineering discovery environment) which is supported by NSF grant numbers ACI-1548562, and (2) ACCESS (Advanced cyberinfrastructure coordination ecosystem: services & support) which is supported by NSF grants numbers #2138259, #2138286, #2138307, #2137603, and #2138296. Specifically, we used the resources from SDSC Expanse GPU compute nodes, and NCSA Delta system, via allocations TG-CIS220009.
@@ Appendix A Glossary
    $\bar{G}_{\ell}\equiv\frac{1}{n}\sum_{\alpha=1}^{n}\frac{\partial f(x^{(\alpha)})}{\partial h_{\ell}}\frac{\partial f(x^{(\alpha)})}{\partial h_{\ell}}^{\top}.$
@@ Appendix B Connection of the PTK to the empirical NTK
Appendix C Additional proofs and statements
Theorem (Maximum C-NFC).
Suppose the data $X$ are sampled uniformly at random from $\mathbb{S}^{d-1}$ and the labels are generated by $f^{*}$ satisfying Assumptions 4 and 5. Suppose we train $f$ with MSE loss and initialize the network so that the initial outputs are 00 by either Method 7 or Method 8 above. Assume the activation function $\phi$ satisfies Assumption 6. We consider the regime that $\omega_{d}\left(d\log^{2\ell+1}d\right)\leq n\leq o_{d}\left(d^{2-\delta}\right)$ for some $\delta>0$, and width $k\rightarrow\infty$. Then, at $t=0$ in training (in the setting of Proposition 3), almost surely over the random $X$ as $n,d\rightarrow\infty$,
$\rho\left(\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}\bar{W}),\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}K\bar{W})\right)\rightarrow 1~{}.$
Proof of Theorem 9.
Applying Proposition 3 and using that we train with MSE and zero initial outputs, we have
$\frac{1}{2}\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}\bar{W})=X^{\top}Y\mathcal{K}YX,\quad\frac{1}{2}\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}\bar{W})=X^{\top}Y\mathcal{K}^{2}YX~{}.$
Given that $k\rightarrow\infty$, $\mathcal{K}$ is equal to a deterministic kernel matrix conditioned on the inputs $X$ (i.e. does not depend on the sampled initial weights). We then use that we are in the proportional limit to approximate the kernel matrix $\mathcal{K}$ by its Taylor expansion (El Karoui, 2008, Hu and Lu, 2022, Ba et al., 2019, Misiakiewicz, 2022). Namely, we have that with probability $1-o_{d}(1)$,
$\mathcal{K}=\gamma\underline{1}\underline{1}^{\top}+\frac{\alpha}{d}XX^{\top}+\mu I+o_{d}(1)\cdot\Delta~{},$ (10)
where $\gamma=h_{d}(0)+\frac{1}{d}h^{\prime\prime}_{d}(0)$, $\alpha=h^{\prime}(0)$, $\mu=h(1)-h(0)-h^{\prime}(0)$, and $\Delta$ with $\|\Delta\|_{2}=1$ is a matrix. In general, we will write standalone asymptotic variables to indicate matrices $\Delta$ of that order spectral norm.
By the linearization in equation (10) and the structure of $XX^{\top}$, we will show,
$\mathcal{K}=\Theta(1)\cdot\underline{1}\underline{1}^{\top}+\widetilde{O}\left(1\right)\cdot\Delta_{1},\quad\mathcal{K}^{2}=\Theta(d)\cdot\underline{1}\underline{1}^{\top}+\widetilde{O}\left(d\right)\cdot\Delta_{2}~{},$
where $\Delta_{1},\Delta_{2}$ have spectral norm $1$. To prove this, first note that $\frac{1}{n}X^{\top}X=I+o(1)$ as uniform data on the sphere are $O(1)$-subgaussian and have identity covariance. Therefore, as $n=\widetilde{\Theta}(d)$, and the spectra of $X^{\top}X$ and $XX^{\top}$ are identical, we have that $\|\frac{1}{d}XX^{\top}\|=\|\frac{1}{d}X^{\top}X\|=\widetilde{O}(1)$.
We now analyze the numerator in the correlation (using the normalized trace). To simplify notation we write $\langle A,B\rangle\equiv\text{tr}\left(A^{\top}B\right)$ for matrices $A,B$. The numerator is then:
$\langle X^{\top}Y\mathcal{K}YX,X^{\top}Y\mathcal{K}^{2}YX\rangle$ $=\widetilde{\Theta}(d)\cdot\text{tr}\left((X^{\top}yy^{\top}X)^{2}\right)$
$+\widetilde{O}(d)\cdot\text{tr}\left((X^{\top}Y\Delta_{1}YX)(X^{\top}yy^{\top}X)\right)$
$+\widetilde{O}(d)\cdot\text{tr}\left((X^{\top}yy^{\top}X)(X^{\top}Y\Delta_{2}YX)\right)$
$+\widetilde{O}(d)\cdot\text{tr}\left((X^{\top}Y\Delta_{1}YX)(X^{\top}Y\Delta_{2}YX)\right)$
As the labels are bounded, $\|Y\Delta_{1}Y\|,\|Y\Delta_{2}Y\|=\widetilde{O}(1)$. Therefore, by the sub-multiplicative property of the spectral norm and that $\|X^{\top}\|\|X\|=\|X^{\top}X\|$, we have $\|X^{\top}Y\Delta_{1}YX\|\leq\|Y\Delta_{1}Y\|\|X^{\top}X\|=\widetilde{O}(1)\cdot\|X^{\top}X\|$ and $\|X^{\top}Y\Delta_{2}YX\|=\widetilde{O}(1)\cdot\|X^{\top}X\|$. Finally, again applying $\|X^{\top}X\|=\widetilde{O}(d)$,
$\text{tr}\left((X^{\top}Y\Delta_{1}YX)(X^{\top}yy^{\top}X)\right)\leq\widetilde{O}(1)\cdot\|X^{\top}X\|\cdot\|X^{\top}y\|^{2}=\widetilde{O}(d)\cdot\|X^{\top}y\|^{2},$
and,
$\text{tr}\left((X^{\top}yy^{\top}X)(X^{\top}Y\Delta_{2}YX)\right)\leq\widetilde{O}(d)\cdot\|X^{\top}y\|^{2}.$
The fourth trace then satisfies,
$\text{tr}\left((X^{\top}Y\Delta_{1}YX)(X^{\top}Y\Delta_{2}YX)\right)\leq\widetilde{O}(1)\cdot\|X^{\top}X\|^{2}=\widetilde{O}(d^{2})$
Combining the four traces and using that $\text{tr}\left((X^{\top}yy^{\top}X)^{2}\right)=\text{tr}\left((y^{\top}XX^{\top}y)^{2}\right)=\|X^{\top}y\|^{4}$,
$\langle X^{\top}Y\mathcal{K}YX,X^{\top}Y\mathcal{K}^{2}YX\rangle$ $=\widetilde{\Omega}(d)\cdot\|X^{\top}y\|^{4}+\widetilde{O}(d^{2})\cdot\|X^{\top}y\|^{2}+\widetilde{O}(d^{3})~{}.$
Recall by Assumption 5, we have $\|\mathbb{E}_{x}\left[xf^{*}(x)\right]\|=\Omega(1)$. We apply Lemma 13 for our choice of $n$ so that $\|\frac{1}{n}X^{\top}y-\mathbb{E}\left[xf^{*}(x)\right]\|=o(1)$ w.p. $1-o(1)$. As a consequence, we have $\|X^{\top}y\|=\widetilde{\Omega}(d)$ w.h.p. Therefore, the leading term in the numerator of the correlation is that of $\text{tr}\left((X^{\top}yy^{\top}X)^{2}\right)$ the term due to the interaction of the mean terms of $\mathcal{K}$ and $\mathcal{K}^{2}$.
Meanwhile the first denominator term decomposes as follows:
$\langle X^{\top}Y\mathcal{K}YX,X^{\top}Y\mathcal{K}YX\rangle$ $=\widetilde{\Theta}(1)\cdot\text{tr}\left((X^{\top}yy^{\top}X)^{2}\right)$
$+\widetilde{O}(1)\cdot\text{tr}\left((X^{\top}Y\Delta_{1}YX)(X^{\top}yy^{\top}X)\right)$
$+\widetilde{O}(1)\cdot\text{tr}\left((X^{\top}Y\Delta_{1}YX)^{2}\right)$
By a similar argument to the numerator term, we observe that the leading sub-term for the first denominator term is $\widetilde{\Theta}(1)\cdot\text{tr}\left((X^{\top}yy^{\top}X)^{2}\right)$.
Further, the second denominator terms decomposes as follows:
$\langle X^{\top}Y\mathcal{K}^{2}YX,X^{\top}Y\mathcal{K}^{2}YX\rangle$ $=\widetilde{\Theta}(d^{2})\cdot\text{tr}\left((X^{\top}yy^{\top}X)^{2}\right)$
$+\widetilde{O}(d^{2})\cdot\text{tr}\left((X^{\top}Y\Delta_{2}YX)(X^{\top}yy^{\top}X)\right)$
$+\widetilde{O}(d^{2})\cdot\text{tr}\left((X^{\top}Y\Delta_{2}YX)^{2}\right)$
By a similar argument, the second denominator term has leading sub-term $\Theta(d^{2})\cdot\text{tr}\left((X^{\top}yy^{\top}X)^{2}\right)$.
Putting the numerator and denominator terms together, we have the correlation has the following asymptotics:
$\rho\left(\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}\bar{W}),\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}K\bar{W})\right)$ $=\frac{\widetilde{\Theta}(d)\cdot\text{tr}\left((X^{\top}yy^{\top}X)^{2}\right)+\widetilde{O}(d^{3})}{\text{tr}\left(\widetilde{\Theta}(1)\cdot(X^{\top}yy^{\top}X)^{2}+\widetilde{O}(d^{2})\right)^{1/2}\cdot\text{tr}\left(\widetilde{\Theta}(d^{2})\cdot(X^{\top}yy^{\top}X)^{2}+\widetilde{O}(d^{4})\right)^{1/2}}$
$\rightarrow 1~{},$
completing the proof. ∎
Lemma 13.
Suppose $|f^{*}(x)|\leq c\cdot\log^{\ell}(d)$ with probability $1-o(1)$ for constant $c>0$, and the data are sub-Gaussian with constant parameter $K$ and some $\ell>0$. Then, $\|\frac{1}{n}X^{\top}y-\mathbb{E}\left[xf^{*}(x)\right]\|=o(1)$ w.p. $1-o(1)$, provided $n=\omega(d\log^{2\ell+1}(d))$.
Proof.
As the data $x$ are $K$-sub-Gaussian for a universal constant $K$, and the labels are bounded by $c\log^{\ell}(d)$, we have that $xy$ is sub-Gaussian with parameter $M=Kc\log^{\ell}(d)$. Therefore, the empirical expectation of $xy$, $\frac{1}{n}\sum_{i=1}^{n}x_{i}y_{i}$ is sub-Gaussian with parameter $M/\sqrt{n}$. Directly applying Lemma 1 in Jin et al. (2019), we see that,
$\mathrm{Pr}\left(\left\|\frac{1}{n}\sum_{i=1}^{n}x_{i}y_{i}-\mathbb{E}\left[xy\right]\right\|>t\right)<2\exp\left(-\frac{t^{2}n}{2cM^{2}d}\right)~{},$
for a universal constant $c>0$. Therefore, provided $n=\omega(d\log^{2\ell+1}(d))$, we have $\left\|\frac{1}{n}\sum_{i=1}^{n}x_{i}y_{i}-\mathbb{E}\left[xy\right]\right\|=o(1)$ with probability $1-o(1)$, completing the proof. ∎
Appendix D Additional centerings of the NFC
@@
Appendix E Extending our theoretical predictions to depth and general activations
@@
Appendix F Free probability calculations of C-NFC
@@
We assume two properties hold in the finite dimensional case we consider, that will hold asymptotically in the infinite dimensional limit.
@@
F.1 Free probability identities
@@ F.2 Numerator term of NFC
F.3 First denominator term of NFC
@@
F.4 Second denominator term of NFC
@@
Appendix G Alignment reversing dataset
The data consists of a mixture of two distributions from which two subsets of the data $X^{(1)}$ and $X^{(2)}$ are sampled from, and is parametrized by a balance parameter $\gamma\in(0,1]$ and two variance parameters $\epsilon_{1},\epsilon_{2}>0$. The subset $X^{(1)}$ which has label $y_{1}=1$ and constitutes a $\gamma$ fraction of the entire dataset, is sampled from a multivariate Gaussian distribution with mean 00 and covariance $\Sigma=\underline{1}\underline{1}^{\top}+\epsilon_{1}\cdot I~{}.$ Then the second subset, $X^{(2)}$, is constructed such that $(X^{(2)})^{\top}X^{(2)}\approx((X^{(1)})^{\top}X^{(1)})^{-2}$, and has labels $y_{2}=0$. Then, for balance parameter $\gamma$ sufficiently small, the AGOP second derivative approximately satisfies,
$\mathbb{E}\left[\dot{W}^{\top}K\dot{W}\right]$ $\sim X^{\top}YXX^{\top}XX^{\top}YX=(X^{(1)})^{\top}X^{(1)}X^{\top}X(X^{(1)})^{\top}X^{(1)}\approx(X^{(1)})^{\top}X^{(1)}(X^{(2)})^{\top}X^{(2)}(X^{(1)})^{\top}X^{(1)}$
$\approx I~{},$
Specifically, we construct $X^{(2)}$ by the following procedure:
@@ Appendix G Alignment reversing dataset
$\mathbb{E}\left[K^{2}_{ij}\right]$ $=\sum_{s}^{k}\sum_{\alpha}^{n}\sum_{p_{1},p_{2}}^{d}\mathbb{E}\left[a_{s}^{4}W_{s,p_{1}}X_{\alpha,p_{1}}X_{\alpha,p_{2}}W_{s,p_{2}}X_{i}^{\top}W_{s}W_{s}^{\top}X_{j}\right]=\sum_{s}^{k}\sum_{\alpha}^{n}\sum_{p_{1},p_{2},q_{1},q_{2}}^{d}$
@@
Appendix H Varying the data distribution
@@
Appendix I Effect of initialization on feature learning
@@
Appendix J Experimental details
@@
Appendix K Additional SLO experiments
@@
Appendix L Additional C-NFC/UC-NFC plots across architectures
@@
Appendix M Additional experiments on real datasets